# Hamiltonian Descent Algorithms for Optimization: Accelerated Rates via Randomized Integration Time

### v1

@@ Abstract
We study the Hamiltonian flow for optimization (HF-opt), which simulates the Hamiltonian dynamics for some integration time and resets the velocity to 00 to decrease the objective function; this is the optimization analogue of the Hamiltonian Monte Carlo algorithm for sampling. For short integration time, HF-opt has the same convergence rates as gradient descent for minimizing strongly and weakly convex functions. We show that by randomizing the integration time in HF-opt, the resulting randomized Hamiltonian flow (RHF) achieves accelerated convergence rates in continuous time, similar to the rates for the accelerated gradient flow. We study a discrete-time implementation of RHF as the randomized Hamiltonian gradient descent (RHGD) algorithm. We prove that RHGD achieves the same accelerated convergence rates as Nesterov’s accelerated gradient descent (AGD) for minimizing smooth strongly and weakly convex functions. We provide numerical experiments to demonstrate that RHGD is competitive with classical accelerated methods such as AGD across all settings and outperforms them in certain regimes.
Qiang Fu qiang.fu@yale.edu Andre Wibisono andre.wibisono@yale.edu
@@ 2.1 Notations
Let $\|\cdot\|:=\|\cdot\|_{2}$ denote the Euclidean norm on the $d$-dimensional Euclidean space $\mathbb{R}^{d}$. A differentiable function $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is $\alpha$-strongly convex if $f(y)\geq f(x)+\langle\nabla f(x),y-x\rangle+\frac{\alpha}{2}\|y-x\|^{2}$ for any $x,y\in\mathbb{R}^{d}$, where $f$ is (weakly) convex if $\alpha=0$. We say $f$ is $\alpha$-gradient-dominated if $\|\nabla f(x)\|^{2}\geq 2\alpha(f(x)-f(x^{*}))$ for any $x\in\mathbb{R}^{d}$ where $x^{*}=\arg\min_{x\in\mathbb{R}^{d}}{f(x)}$ is a minimizer of $f$. We say $f$ is $L$-smooth if $f(y)\leq f(x)+\langle\nabla f(x),y-x\rangle+\frac{L}{2}\|y-x\|^{2}$ or equivalently $\|\nabla f(x)-\nabla f(y)\|\leq L\|x-y\|$ for any $x,y\in\mathbb{R}^{d}$. We assume $\alpha\leq 1\leq L$. We say $\hat{x}$ is an $\varepsilon$-accurate solution in expectation if $\mathbb{E}[f(\hat{x})-f(x^{*})]\leq\varepsilon$. Let $\kappa:=L/\alpha$ denote the condition number. We use $\dot{g}_{t}:=\frac{\mathrm{d}}{\mathrm{d}t}g_{t}$ to denote the time derivative of a time-dependent quantity $g_{t}$. We define the flow map $\mathsf{HF}_{\eta}:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}\times\mathbb{R}^{d}$ as $\mathsf{HF}_{\eta}(X_{0},Y_{0})=(X_{\eta},Y_{\eta})$, which is the solution to Hamiltonian flow (HF) at time $\eta$ starting from $(X_{0},Y_{0})$. Given a time-dependent random variable $Z_{t}$, we use $\rho_{t}^{Z}$ to denote its probability distribution. We identify probability distributions with their density functions. Let $\mathrm{Exp}(\gamma)$ denote the exponential distribution with mean $1/\gamma$ for $\gamma>0$. Let $[n]:=\{1,2,...,n\}$. We use $a=O(b)$ to denote $a\leq Cb$ for constant $C>0$ and use $a=\tilde{O}(b)$ to denote $a=O(b)$ up to logarithmic factors. We use $a=\Theta(b)$ to denote $a=Cb$ for constants $C>0$.
@@ 2.2 Hamiltonian flow for optimization
We can exploit the conservation property of the Hamiltonian flow to design an optimization algorithm by periodically refreshing the velocity to 00. This idea results in the following Hamiltonian flow for optimization (HF-opt) algorithm, which was also proposed and studied by Teel et al., (2019); Wang, (2024). Below, $\Pi_{1}(x,y)=x$ is the projection operator to the first component.
@@
3:     $x_{k+1}=\Pi_{1}\circ\mathsf{HF}_{\eta_{k}}(x_{k},0)$triangleright (evolve (HF) for time $\eta_{k}$ and project to first component)
@@ 3 Randomized Hamiltonian flow for optimization
5:     Set $x_{k+1}=\Pi_{1}\circ\mathsf{HF}_{\tau_{k}}(x_{k},0)$triangleright (evolve (HF) for time $\tau_{k}$ and project to first component)
@@
[…]
@@ 3.1 Reformulation of RHF-opt as a continuous-time process
    At random jump times governed by an inhomogeneous Poisson process with rate $\gamma(t)$, we refresh the velocity to 00, and continue evolving (HF).
@@ 4.1.2 For weakly convex functions
Assume $f$ is weakly convex and $L$-smooth. Then for all $k\geq 0$, RHGD (Algorithm 3) with $h\leq\frac{1}{8\sqrt{L}}$, $\gamma_{k}=\frac{17}{2(k+9)h}$, and from any $x_{0}\in\mathbb{R}^{d}$ satisfies $\mathbb{E}[f(x_{k})-f(x^{*})]\leq\frac{14\cdot\mathbb{E}\left[\|x_{0}-x^{*}\|^{2}\right]}{h^{2}(k+8)^{2}}.$
@@
Assume $f$ is weakly convex and $L$-smooth. To generate $x_{K}$ satisfying $\mathbb{E}[f(x_{K})-f(x^{*})]\leq\varepsilon$, it suffices to run Algorithm 3 with $h=\frac{1}{8\sqrt{L}}$, $\gamma_{k}=\frac{17}{2(k+9)h}$ and any $x_{0}\in\mathbb{R}^{d}$ for
$K\geq\sqrt{\frac{896L\cdot\mathbb{E}\left[\|x_{0}-x^{*}\|^{2}\right]}{\varepsilon}}.$
@@ 5.1 Minimizing quadratic functions
We consider both strongly convex ($\alpha>0$) and weakly convex ($\alpha=0$) quadratic minimization. For the strongly convex case, we fix $L=500$ and test three condition numbers $\kappa\in\{10^{3},10^{5},10^{7}\}$ with $\alpha=L/\kappa$. Since AGD, CAGD, and RHGD require $\alpha$ to compute momentum parameters and refreshment rates, we first compare all algorithms using the exact value of $\alpha$. Notably, when the condition number $\kappa$ is large, the strong convexity constant $\alpha$ can be extremely small (e.g., $\kappa=10^{7}$, $\alpha=5\times 10^{-5}$), which makes an accurate estimation of $\alpha$ challenging in practice. To evaluate robustness, we also test with misspecified values $\hat{\alpha}\in\{0.01,0.1,1\}>\alpha$. In this section, we report results for $\kappa=10^{7}$ and $\hat{\alpha}=0.01$; additional results are deferred to Appendix H.1.1. For the weakly convex case, we evaluate the same algorithms with appropriately configured parameters (see Appendix H.3). In particular, RHGD uses decaying $\gamma_{k}=\frac{17}{2(k+9)h}$ instead of a constant.
Summary of experimental results. Figure LABEL:fig:quadcomp presents results for the strongly convex setting with exact $\alpha$ (left), misspecified $\hat{\alpha}=0.01$ (middle), and the weakly convex setting (right). In the left plot, all accelerated methods clearly outperform GD. While slower than AGD and CAGD, which are known to be optimal under exact parameter knowledge, RHGD remains highly competitive during the early iterations. The middle plot highlights the robustness of RHGD to misspecification. The momentum parameters of AGD and CAGD, and the refreshment rate of RHGD are computed using overestimated values $\hat{\alpha}=0.01$, which are significantly larger than the true value $\alpha=5\times 10^{-5}$. While AGD and CAGD degrade significantly when using overestimated $\hat{\alpha}$, RHGD maintains stable performance and outperforms them in later stages. The right plot shows that RHGD achieves the fastest convergence among all algorithms, consistently improving over AGD and CAGD throughout the iterations, demonstrating its advantage in the weakly convex setting.
@@ 5.2 Minimizing logistic regression loss
[…]
@@ Appendix A Related work
[…]
@@ Appendix B Lift-Conserve-Project scheme
  • A conservation map $\bf{C}\colon\mathcal{Z}\rightarrow\cal{Z}$ which preserves $\cal{H}$, i.e., $\mathcal{H}(z)=\mathcal{H}(\mathbf{C}(z)).$
@@ B.2.1 Distributional properties of Hamiltonian flow
Let $(x_{t},y_{t})=\textsf{\em HF}_{t}(x_{0},y_{0})$ denote the solution to (HF) at time $t$ where $(x_{t},y_{t})\sim\rho_{t}$. Then we have
$\mathsf{KL}(\rho_{t}\,\|\,\tilde{\nu})=\mathsf{KL}(\rho_{0}\,\|\,\tilde{\nu}).$ (7)
@@
Computing the time derivative of $\mathsf{KL}(\rho_{t}\|\tilde{\nu})$ along (HF), we obtain
$\frac{\mathrm{d}}{\mathrm{d}t}\mathsf{KL}(\rho_{t}\|\tilde{\nu})$ $=\frac{\mathrm{d}}{\mathrm{d}t}\int\rho_{t}\log\frac{\rho_{t}}{\tilde{\nu}_{t}}=\int\partial_{t}\rho_{t}\log\frac{\rho_{t}}{\tilde{\nu}_{t}}+\int\rho_{t}\frac{\partial_{t}\rho_{t}}{\rho_{t}}$
=(7)${=}-\int\nabla\cdot\left(\rho_{t}\left(\begin{array}[]{c}y\\ -\nabla f(x)\end{array}\right)\right)\log\frac{\rho_{t}}{\tilde{\nu}_{t}}+\underbrace{\int\partial_{t}\rho_{t}}_{=0}$
@@
where $\nabla:=(\nabla_{x},\nabla_{y})^{\top}$, and the third and the last equations follow from integration by part. Thus $\frac{\mathrm{d}}{\mathrm{d}t}\mathsf{KL}(\rho_{t}\|\tilde{\nu})=0$, which implies $\mathsf{KL}(\rho_{t}\|\tilde{\nu})=\mathsf{KL}(\rho_{0}\|\tilde{\nu}).$
@@ B.2.2 Hamiltonian Monte Carlo as an LCP scheme
In the initial lifting, since $\rho_{0}^{Y\mid X}=\tilde{\nu}^{Y\mid X}=\mathcal{N}(0,I)$, by Lemma 5 we have $\textsf{KL}(\rho_{0}^{X}\,\|\,\nu)=\textsf{KL}(\rho_{0}\,\|\,\tilde{\nu})$. Then we have
@@ C.1 Review of optimization algorithms with convergence rates
[…]
@@ C.1.1 Convergence under strong convexity or gradient domination
We can show ˙$\dot{\mathcal{E}}_{t}\leq 0$ along (AGF) with $\beta(t)=2\sqrt{\alpha}$ for $t\geq 0$. For a complete proof, see the proof of Wilson et al., (2021, Proposition 3). ∎
@@ C.1.2 Convergence under weak convexity
Define the continuous-time Lyapunov function: $\mathcal{E}_{t}=\frac{1}{2}\|X_{t}-x^{*}\|^{2}+t(f(X_{t})-f(x^{*}))$. We can show ˙$\dot{\mathcal{E}}_{t}\leq 0$ along (GF) for $t\geq 0$. See Wilson, (2018, Appendix B.2.2) for more details. ∎
@@
We can show ˙$\dot{\mathcal{E}}_{t}\leq 0$ along (AGF) with $\beta(t)=3/t$ for $t\geq 0$. A complete proof can be found in the proof of Su et al., (2016, Theorem 3). ∎

### v2

@@ Abstract
We study the Hamiltonian flow for optimization (HF-opt), which simulates the Hamiltonian dynamics for some integration time and resets the velocity to 0 to decrease the objective function; this is the optimization analogue of the Hamiltonian Monte Carlo algorithm for sampling. For short integration time, HF-opt has the same convergence rates as gradient descent for minimizing strongly and weakly convex functions. We show that by randomizing the integration time in HF-opt, the resulting randomized Hamiltonian flow (RHF) achieves accelerated convergence rates in continuous time, similar to the rates for the accelerated gradient flow. We study a discrete-time implementation of RHF as the randomized Hamiltonian gradient descent (RHGD) algorithm. We prove that RHGD achieves the same accelerated convergence rates as Nesterov’s accelerated gradient descent (AGD) for minimizing smooth strongly and weakly convex functions. We provide numerical experiments to demonstrate that RHGD is competitive with classical accelerated methods such as AGD across all settings and outperforms them in certain regimes.
Qiang Fu
qiang.fu@yale.edu
Andre Wibisono
andre.wibisono@yale.edu
@@ 2.1 Notations
Let $\|\cdot\|:=\|\cdot\|_{2}$ denote the Euclidean norm on the $d$-dimensional Euclidean space $\mathbb{R}^{d}$. A differentiable function $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is $\alpha$-strongly convex if $f(y)\geq f(x)+\langle\nabla f(x),y-x\rangle+\frac{\alpha}{2}\|y-x\|^{2}$ for any $x,y\in\mathbb{R}^{d}$, where $f$ is (weakly) convex if $\alpha=0$. We say $f$ is $\alpha$-gradient-dominated if $\|\nabla f(x)\|^{2}\geq 2\alpha(f(x)-f(x^{*}))$ for any $x\in\mathbb{R}^{d}$ where $x^{*}=\arg\min_{x\in\mathbb{R}^{d}}{f(x)}$ is a minimizer of $f$. We say $f$ is $L$-smooth if $f(y)\leq f(x)+\langle\nabla f(x),y-x\rangle+\frac{L}{2}\|y-x\|^{2}$ or equivalently $\|\nabla f(x)-\nabla f(y)\|\leq L\|x-y\|$ for any $x,y\in\mathbb{R}^{d}$. We assume $\alpha\leq 1\leq L$. We say $\hat{x}$ is an $\varepsilon$-accurate solution in expectation if $\mathbb{E}[f(\hat{x})-f(x^{*})]\leq\varepsilon$. Let $\kappa:=L/\alpha$ denote the condition number. We use $\dot{g}_{t}:=\frac{\mathrm{d}}{\mathrm{d}t}g_{t}$ to denote the time derivative of a time-dependent quantity $g_{t}$. We define the flow map $\mathsf{HF}_{\eta}:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}\times\mathbb{R}^{d}$ as $\mathsf{HF}_{\eta}(X_{0},Y_{0})=(X_{\eta},Y_{\eta})$, which is the solution to Hamiltonian flow (HF) at time $\eta$ starting from $(X_{0},Y_{0})$. Given a time-dependent random variable $Z_{t}$, we use $\rho_{t}^{Z}$ to denote its probability distribution. We identify probability distributions with their density functions. Let $\mathrm{Exp}(\gamma)$ denote the exponential distribution with mean $1/\gamma$ for $\gamma>0$. Let $n]:=\{1,2,...,n\}$. We use $a=O(b)$ to denote $a\leq Cb$ for constant $C>0$ and use $a=\tilde{O}(b)$ to denote $a=O(b)$ up to logarithmic factors. We use $a=\Theta(b)$ to denote $a=Cb$ for constants $C>0$.
@@ 2.2 Hamiltonian flow for optimization
We can exploit the conservation property of the Hamiltonian flow to design an optimization algorithm by periodically refreshing the velocity to 0. This idea results in the following Hamiltonian flow for optimization (HF-opt) algorithm, which was also proposed and studied by Teel et al., (2019); Wang, (2024). Below, $\Pi_{1}(x,y)=x$ is the projection operator to the first component.
@@
3:  $x_{k+1}=\Pi_{1}\circ\mathsf{HF}_{\eta_{k}}(x_{k},0)$ $\triangleright$ (evolve (HF) for time $\eta_{k}$ and project to first component)
@@ 3 Randomized Hamiltonian flow for optimization
5:  Set $x_{k+1}=\Pi_{1}\circ\mathsf{HF}_{\tau_{k}}(x_{k},0)$ $\triangleright$ (evolve (HF) for time $\tau_{k}$ and project to first component)
@@
In Algorithm 2, the $k$-th integration time $\tau_{k}$ is a random variable drawn from an exponential distribution with mean $1/\gamma(T_{k})$, where $T_{k}$ is the current time. Note that $\gamma(t)$ can depend on time. Later in Section 3.2, we choose $\gamma(t)$ to be a constant when $f$ is strongly convex, and $\gamma(t)\,\propto\,1/t$ when $f$ is weakly convex. We first illustrate the behavior of RHF-opt on a simple quadratic function. See Appendix E for the proof.
@@ 3.1 Reformulation of RHF-opt as a continuous-time process
    At random jump times governed by an inhomogeneous Poisson process with rate $\gamma(t)$, we refresh the velocity to 0, and continue evolving (HF).
@@ 4.1.2 For weakly convex functions
Assume $f$ is weakly convex and $L$-smooth. Then for all $k\geq 0$, RHGD (Algorithm 3) with $h\leq\frac{1}{7\sqrt{L}}$, $\gamma_{k}=\frac{17}{2(k+9)h}$, and from any $x_{0}\in\mathbb{R}^{d}$ satisfies $\mathbb{E}[f(x_{k})-f(x^{*})]\leq\frac{14\cdot\mathbb{E}\left[\|x_{0}-x^{*}\|^{2}\right]}{h^{2}(k+8)^{2}}.$
@@
Assume $f$ is weakly convex and $L$-smooth. To generate $x_{K}$ satisfying $\mathbb{E}[f(x_{K})-f(x^{*})]\leq\varepsilon$, it suffices to run Algorithm 3 with $h=\frac{1}{7\sqrt{L}}$, $\gamma_{k}=\frac{17}{2(k+9)h}$ and any $x_{0}\in\mathbb{R}^{d}$ for
$K\geq\sqrt{\frac{686L\cdot\mathbb{E}\left[\|x_{0}-x^{*}\|^{2}\right]}{\varepsilon}}.$
@@ 5.1 Minimizing quadratic functions
[…]
(a) Comparison between GD, AGD, CAGD, and RHGD (ours) on minimizing quadratic functions with $L=500$ in three settings: (1) $\kappa=10^{7}$ with exact $\alpha$ (left); (2) $\kappa=10^{7}$ with misspecified $\hat{\alpha}=0.01$ (middle); (3) $\alpha=0$ (right). We use optimal stepsizes via grid search for each setting. Each plot shows results averaged over 5 runs.
Summary of experimental results. Figure 3(a) presents results for the strongly convex setting with exact $\alpha$ (left), misspecified $\hat{\alpha}=0.01$ (middle), and the weakly convex setting (right). In the left plot, all accelerated methods clearly outperform GD. While slower than AGD and CAGD, which are known to be optimal under exact parameter knowledge, RHGD remains highly competitive during the early iterations. The middle plot highlights the robustness of RHGD to misspecification. The momentum parameters of AGD and CAGD, and the refreshment rate of RHGD are computed using overestimated values $\hat{\alpha}=0.01$, which are significantly larger than the true value $\alpha=5\times 10^{-5}$. While AGD and CAGD degrade significantly when using overestimated $\hat{\alpha}$, RHGD maintains stable performance and outperforms them in later stages. The right plot shows that RHGD achieves the fastest convergence among all algorithms, consistently improving over AGD and CAGD throughout the iterations, demonstrating its advantage in the weakly convex setting.
@@ 5.2 Minimizing logistic regression loss
(b) Comparison of GD, AGD, CAGD and RHGD on logistic regression with $\alpha\in\{10^{-4},0\}$. We run each algorithm using adaptive stepsizes. For RHGD, we evaluate $\gamma_{k}=\sqrt{\alpha}$ and $\gamma_{k}=2\sqrt{\alpha}$ for $\alpha>0$, and evaluate $\gamma_{k}=17/(2k+18)h$ for $\alpha=0$.
Summary of experimental results. Figure 3(b) presents the convergence behavior of all tested algorithms on a logistic regression task with $\alpha\in\{10^{-4},\,0\}$ and adaptive stepsizes. We observe from the left plot that using a slightly larger refreshment rate, $\gamma_{k}=2\sqrt{\alpha}$, results in faster and more stable convergence, while Theorem 4 suggests setting the refreshment rate to $\gamma_{k}=\sqrt{\alpha}$. Our proposed algorithm RHGD with $\gamma_{k}=2\sqrt{\alpha}$ and $\gamma_{k}=\sqrt{\alpha}$ are faster than GD in the late iterations and comparable to AGD and CAGD. The right plot presents the results in the weakly convex setting with adaptive stepsizes. For RHGD, we choose the decaying refreshment rate $\gamma_{k}=\frac{17}{2(k+9)h}$. All accelerated methods consistently outperform GD, and RHGD exhibits comparable performance to baseline accelerated algorithms AGD and CAGD.
@@ Appendix A Related work
To the best of our knowledge, only a few previous works have explored optimization algorithms explicitly based on energy-conserving Hamiltonian flow of the form (HF). Teel et al., (2019) studied optimization methods derived from Hamiltonian flow with a velocity refreshment strategy. Specifically, they refresh the velocity to zero whenever the iterate approaches the boundary of a region defined as $\{(x,y)\in\mathbb{R}^{d}\times\mathbb{R}^{d}:\langle\nabla f(x),y\rangle\leq 0,\|y\|^{2}\geq\|\nabla f(x)\|^{2}/L\}$ where $L$ is the smoothness constant, or after a fixed timer expires. They established uniform global stability and convergence guarantees for minimizing smooth and strongly convex functions. Diakonikolas and Jordan, (2021) analyzed generalized Hamiltonian dynamics characterized by a time-dependent Hamiltonian $H(x,y,\tau)=h(\tau)f(x/\tau)+\psi^{*}(y),$ where $h(\tau)$ is a positive function of scaled time $\tau$ and $\psi^{*}(\cdot)$ is strongly convex and differentiable. They showed that along these Hamiltonian flows, the average gradient norm $\|\frac{1}{t}\int_{0}^{t}\nabla f(x_{\tau})d\tau\|$ decreases at an O(1/t)O({1}/{t}) rate. Moreover, they demonstrated that a broad family of momentum methods, applicable in both Euclidean and non-Euclidean spaces, can be derived from these generalized dynamics. This family includes classical methods such as Nesterov’s accelerated gradient method (Nesterov,, 1983, 2018), Heavy-Ball method (Polyak,, 1964), and other well-known accelerated methods (Wibisono et al.,, 2016; Wilson et al.,, 2021) as special cases, which are different from what we consider in this paper. De Luca et al., (2023) propose a Hamiltonian-based method with the Hamiltonian defined as $H(x,y)=\lambda\log(F(x)-F_{0})+\log(\|y\|^{2})$ where $\lambda>0$ and $F_{0}$ are user-specified parameters. They show that the proposed method together with some heuristics is competitive with Adam (Kingma and Ba,, 2015) in some deep learning experiments, but no convergence-rate analysis was performed in the work. Wang, (2024) studies the Hamiltonian flow with velocity refreshment for minimizing strongly convex quadratic functions $f(x)=\frac{1}{2}x^{\top}Ax-b^{\top}x$ with $A\in\mathbb{R}^{d\times d}$ and $\alpha I\preceq A\preceq LI$; importantly, (HF) can be solved exactly when $f$ is quadratic, which facilitates the analysis. By choosing the integration time $\eta_{k}$ related to the roots of Chebyshev polynomials, Wang, (2024) shows that HF-opt achieves the convergence rate of $O((1-\Theta(1/\sqrt{\kappa}))^{k})$ and the total integration time ${\Theta}({1}/{\sqrt{\alpha}})$, which matches the total integration of accelerated gradient flow with refreshment (Suh et al.,, 2022).
@@ Appendix B Lift-Conserve-Project scheme
[…]
@@ B.2.1 Distributional properties of Hamiltonian flow
Let $(x_{t},y_{t})=\textsf{HF}_{t}(x_{0},y_{0})$ denote the solution to (HF) at time $t$ where $(x_{t},y_{t})\sim\rho_{t}$. Then we have
[…]
@@
[…]
@@
[…]
@@ B.2.2 Hamiltonian Monte Carlo as an LCP scheme
[…]
@@ C.1 Review of optimization algorithms with convergence rates
$\left\{\begin{aligned} x_{k+1}&=y_{k}-\eta\nabla f(y_{k}),\\ y_{k+1}&=x_{k+1}+\beta_{k}(x_{k+1}-x_{k}),\end{aligned}\right$. (AGD)
@@ C.1.1 Convergence under strong convexity or gradient domination
We can show $\dot{\mathcal{E}}_{t}\leq 0$ along (AGF) with $\beta(t)=2\sqrt{\alpha}$ for $t\geq 0$. For a complete proof, see the proof of Wilson et al., (2021, Proposition 3). ∎
@@ C.1.2 Convergence under weak convexity
Define the continuous-time Lyapunov function: $\mathcal{E}_{t}=\frac{1}{2}\|X_{t}-x^{*}\|^{2}+t(f(X_{t})-f(x^{*}))$. We can show $\dot{\mathcal{E}}_{t}\leq 0$ along (GF) for $t\geq 0$. See Wilson, (2018, Appendix B.2.2) for more details. ∎
@@
We can show $\dot{\mathcal{E}}_{t}\leq 0$ along (AGF) with $\beta(t)=3/t$ for $t\geq 0$. A complete proof can be found in the proof of Su et al., (2016, Theorem 3). ∎